What is the rms value of current i=3+4sin(wt+pi\3)?


Please give step by step solution.

Grade:12th Pass

1 Answers

Sher Mohammad IIT Delhi
9 years ago
y=a\sin(2\pi ft)\,, rms value is\frac{a}{\sqrt{2}}

ans: 3+4/sqrt(2)

sher mohammad
askiitians faculty
b.tech, iit delhi